# Knowledge Independence Breeds Disruption but Limits Recognition

### v1

@@ Acknowledgments
Table S4: Normalized mediation regression on team freshness. We perform normalized mediation analysis to assess the effect of team freshness on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, and team distance, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team freshness, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team freshness KI 0.0007*** -$0**$0004***
@@
Team freshness Novelty $0**$0004*** -0.0004***
@@
Team freshness Reference count -0.0001*** -0.0003*** -0.0004***
@@
Team freshness Reference age -$0**$0003*** -0.0004***
@@
Team freshness Reference $C_{5}$$0**$0004*** -0.0004***
@@
Team freshness Team size -$0**$0004*** -0.0004***
@@
Team freshness Team distance -0.0003*** -0.0001*** -0.0004***
@@
Table S5: Normalized mediation regression on paper impact. We perform normalized mediation analysis to assess the effect of paper impact on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C5$, team size, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by paper impact, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Citation percentile KI -0.0989*** -$0**$0989***
@@
Citation percentile Novelty 0.0003*** -0.0991*** -0.0989***
@@
Citation percentile Reference count -0.0438*** -0.0551*** -0.0989***
@@
Citation percentile Reference age -0.0091*** -0.0897*** -0.0989***
@@
Citation percentile Reference $C_{5}$0.0007*** -0.0982*** -0.0989***
@@
Citation percentile Team size -0.0003*** -0.0986*** -0.0989***
@@
Citation percentile Team distance -0.0041*** -0.0948*** -0.0989***
@@
Citation percentile Team freshness $0**$0989*** -0.0989***
@@
Table S6: Normalized regression models of disruption and the influencing factors in SciSciNet dataset. We perform normalized ordinary-least-squares (OLS) regression analyses to examine the effect of KI on disruption in SciSciNet dataset, while controlling for various confounding factors. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 Model 9 Model 10 Model 11
KI 0.3962*** 0.3831*** 0.3952*** 0.3766*** 0.3984*** 0.3960*** 0.2636*** 0.3959*** 0.3950*** 0.3925*** 0.2300***
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
Log $C_{5}$0.0469*** -0.0497***
@@
Reference disruption 0.2237*** 0.2334***
@@
Team size -0.0096*** -0.0078***
@@
Team distance -0.0235*** -0.0149***
@@
Team freshness 0.0382*** 0.0329***
@@
Fixed discipline Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Fixed year Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
N 19916656 19916656 19916656 19916656 19916656 19916656 19916656 19916656 19916656 19916656 19916656
$R^{2}$ 0.198 0.200 0.199 0.201 0.199 0.199 0.223 0.199 0.199 0.200 0.231
Table S7: Normalized mediation regression on reference disruption in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of reference disruption on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by reference disruption, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Reference disruption KI 0.1729*** 0.2237*** 0.3966***
@@
Reference disruption Novelty -0.0000*** 0.3966*** 0.3966***
@@
Reference disruption Reference count 0.0231*** 0.3735*** 0.3966***
@@
Reference disruption Reference age -0.0037*** 0.4003*** 0.3966***
@@
Reference disruption Reference $C_{5}$0.0003*** 0.3969*** 0.3966***
@@
Reference disruption Team size 0.0003*** 0.3963*** 0.3966***
@@
Reference disruption Team distance 0.0014*** 0.3952*** 0.3966***
@@
Reference disruption Team freshness 0.0037*** 0.3929*** 0.3966***
@@
Table S8: Normalized mediation regression on team size in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of team size on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team size, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team size KI -0.0106*** -0.0096*** -0.0203***
@@
Team size Novelty -0.0005*** -0.0198*** -0.0203***
@@
Team size Reference count -0.0007*** -0.0196*** -0.0203***
@@
Team size Reference age -0.0033*** -0.0169*** -0.0203***
@@
Team size Reference $C_{5}$0.0005*** -0.0198 -0.0203***
@@
Team size Team distance -0.0059*** -0.0144*** -0.0203***
@@
Team size Team freshness 0.0052*** -0.0255*** -0.0203***
@@
Table S9: Normalized mediation regression on team distance in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of team distance on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team distance, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team distance KI -0.0193*** -0.0235*** -0.0429***
@@
Team distance Novelty -0.0003*** -0.0425*** -0.0429***
@@
Team distance Reference count -$0**$0319*** -0.0429***
@@
Team distance Reference age -$0**$0419*** -0.0429***
@@
Team distance Reference $C_{5}$0.0002*** -0.0426*** -0.0429***
@@
Team distance Team size -$0**$0408*** -0.0429***
@@
Table S10: Normalized mediation regression on team freshness in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of team freshness on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, and team distance, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team freshness, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team freshness KI 0.0342*** 0.0382*** 0.0725***
(0.0002) (0.0004) (0.0005)
Team freshness Novelty 0.0001*** 0.0724*** 0.0725***
(0.0000) (0.0005) (0.0005)
Team freshness Reference count 0.0097*** 0.0628*** 0.0725***
(0.0001) (0.0005) (0.0005)
Team freshness Reference age 0.0010*** 0.0715*** 0.0725***
(0.0000) (0.0005) (0.0005)
Team freshness Reference $C_{5}$0.0000*** 0.0725*** 0.0725***
(0.0000) (0.0005) (0.0005)
Team freshness Team size -0.0017*** 0.0742*** 0.0725***
(0.0004) (0.0006) (0.0005)
Team freshness Team distance -0.0013*** 0.0737*** 0.0725***
(0.0000) (0.0005) (0.0005)

### v2

@@ Acknowledgements
Team size Novelty -0.0004*** -0.0212*** -0.0217*** 0.0204
@@
Team size Interdisciplinarity -0.0003*** -0.0214*** -0.0217*** 0.0123
@@
Team size Reference count -0.0006*** -0.0210*** -0.0217*** 0.0289
@@
Team size Reference age -0.0033*** -0.0184*** -0.0217*** 0.1510
@@
Team size Team distance -0.0059*** -0.0158*** -0.0217*** 0.2733
@@
Team size Team freshness 0.0051*** -0.0268*** -0.0217*** 0.2352
@@
  • $\dagger$
    We perform normalized mediation analysis to assess the effect of team size on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, interdisciplinarity, reference count, reference age, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team size, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S8: $|$ Normalized mediation regression on team distance in SciSciNet dataset.
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$) ||ACME / TE||
Team distance KI -0.0203*** -0.0236*** -0.0440*** 0.4630
@@
Team distance Novelty -0.0003*** -0.0437*** -0.0440*** 0.0068
@@
Team distance Interdisciplinarity 0.0007*** -0.0446*** -0.0440*** 0.0152
@@
Team distance Reference count -0.0111*** -0.0329*** -0.0440*** 0.2524
@@
Team distance Reference age -0.0009*** -0.0431*** -0.0440*** 0.0203
@@
Team distance Team size -0.0022*** -0.0417*** -0.0440*** 0.0508
@@
Team distance Team freshness 0.0020*** -0.0459*** -0.0440*** 0.0452
@@
  • $\dagger$
    We perform normalized mediation analysis to assess the effect of team distance on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, interdisciplinarity, reference count, reference age, team size, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team distance, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S9: $|$ Normalized mediation regression on team freshness in SciSciNet dataset.
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$) ||ACME / TE||
Team freshness KI 0.0366*** 0.0393*** 0.0758*** 0.4821
@@
Team freshness Novelty 0.0002*** 0.0757*** 0.0758*** 0.0020
@@
Team freshness Interdisciplinarity 0.0002*** 0.0757*** 0.0758*** 0.0020
(0.0000) (0.0004) (0.0004)
Team freshness Reference count 0.0111*** 0.0648*** 0.0758*** 0.1461
(0.0001) (0.0004) (0.0004)
Team freshness Reference age 0.0011*** 0.0747*** 0.0758*** 0.0150
(0.0000) (0.0004) (0.0004)
Team freshness Team size -0.0018*** 0.0776*** 0.0758*** 0.0232
(0.0004) (0.0006) (0.0004)
Team freshness Team distance -0.0012*** 0.0770*** 0.0758*** 0.0156
(0.0000) (0.0004) (0.0004)
  • $\dagger$
    We perform normalized mediation analysis to assess the effect of team freshness on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, interdisciplinarity, reference count, reference age, team size, and team distance, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team freshness, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S10: $|$ Normalized confounding regression on paper impact in SciSciNet dataset.
Independent variable Confounder Unadjusted ($\alpha_{1}$) Adjusted ($\delta_{1}$) RCE ($|\alpha_{1}-\delta_{1}|/|\alpha_{1}|$)
Citation percentile KI -0.1094*** -0.0184*** 0.8316
@@
Citation percentile Team size -0.1094*** -0.1083*** 0.0097
@@
Citation percentile Team distance -0.1094*** -0.1064*** 0.0270
@@
Citation percentile Team freshness -0.1094*** -0.1053*** 0.0379
@@
  • $\dagger$
    We perform normalized confounding analysis to assess the effect of paper impact on disruption in SciSciNet dataset. The candidate confounding variables include: KI, novelty, interdisciplinarity, reference count, reference age, team size, team distance, and team freshness, with discipline and publication year controlled. The relative change in estimate (RCE) refers to the relative change of effect after controlling for the confounder. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
S7 Reproduced main results based on Web of Science publication dataset
Figure S18: $|$ The distribution and evolution pattern of knowledge independence in Web of Science dataset. (B) The KI distribution of 30,560,343 papers with at least two references published between 1955 and 2017 in the Web of Science. (C) The evolving downtrend of KI over time across disciplines. Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S19: $|$ Knowledge independence is associated with scientific disruption in Web of Science dataset. (A) For the 27,769,300 papers published before 2016 in the Web of Science that are cited at least once, the average disruption percentile (blue curve, left y-axis) and the disruption positive ratio (green curve, right y-axis) increase with the KI percentile. (B) The association between KI and disruption persists, regardless of the impact of the focal paper. Here, impact is measured as the number of citations received within the first five years after publication, denoted by $C_{5}$. (C) The association between KI and disruption persists across decades. (D) The association between KI and disruption persists, regardless of discipline. Bootstrapped 95% confidence intervals are shown as shaded bands in panels A-D. (E) The ATT matrixs of KI on disruption via CEM (upper panel) and PSM (lower panel). Each controlled group is set as a baseline, and ATTs are calculated for comparisons between the baseline and each of the treated groups. Blue cells represent negative ATTs, while red ones represent positive ATTs, with color intensity proportional to the absolute value. Each ATT is tested against the null hypothesis (ATT equals 0) using a two-sided t-test. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Figure S20: $|$ Team’s preference for knowledge independence in Web of Science dataset. (A-C) The KI percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Curves with larger bounds are displayed in the inset to improve visualization. Over time, the relative KI percentile decreases with team size (panel A) and team distance (panel B), and increases with team freshness (panel C); the only exception is observed for the 1950s in the team distance analysis (inset of panel B). (D-F) Solid lines depict the relationships between disruption and team properties with fixed KI values, alongside the uncontrolled cases (insets). Dashed lines represent linear fitted curves. Teams with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are flatter compared to the uncontrolled curves (insets). Bootstrapped 95% confidence intervals are shown as shaded bands. Considering the lack of coordinate information in Web of Science dataset, we categorize the team distance of a given focal paper by hierarchical categorization of the affiliations: 0 (Sub-organization / Department level): 11: Organization / University level; 22: City level; 33: State / Province level; 44: National level; 55: International level. This methodology is also applied in Fig. S25.
Figure S21: $|$ Knowledge independence explains the negative relationship between disruption and impact in Web of Science dataset. We observe a universal negative relationship between papers’ disruption and impact, regardless of the journal rank (panel A), time period (panel B), and discipline (panel C). Here, a paper’s impact is measured by the percentile of its total citations within the same publication year and discipline, and the journal rank is based on Scimago Journal Rank. (D) For papers published in the year $T=2000$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline (see figure S26 for different values of $T$). (E) The citation percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Over time, the relative citation percentile decreases with KI percentile. (F) A paper with a higher KI is more likely to become a Sleeping Beauty (SB coefficient, blue curve, left y-axis) and takes a longer silence to usher in the citation burst (Awaking time, green curve, right y-axis). (G) Solid lines depict the relationships between disruption and impact with fixed KI values, alongside the uncontrolled case (inset). Dashed lines represent linear fitted curves. For papers of comparable impact, those with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are significantly flatter compared to the uncontrolled curve (inset). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S22: $|$ The association between disruption and KI is robust across different reference counts in Web of Science dataset. (A-B) When controlling for the percentile of reference count, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference count. (C-D) When controlling for the reference count within the bottom group (<=10<=10), where each reference count induces a limited range of KI values, both the percentile and positive ratio of disruption continue to increase with KI across all groups of reference count.
Figure S23: $|$ The association between disruption and KI is robust across alternative measures in Web of Science dataset. (A-D) Upper panel: from the perspective of percentile, all disruption measures show a consistent increasing trend with all KI measures. Lower panel: from the perspective of positive ratio, all disruption measures similarly increase with all KI measures. The positive ratio represents the likelihood of having a higher proportion of indind-type references compared to depdep-type references. Consequently, the variation curves for $D_{1}$ and $D_{3}$ align with those for $D_{0}$ and $D_{2}$, respectively.
Figure S24: $|$ The association between disruption and KI is robust when controlling for focal paper’s properties in Web of Science dataset. (A-E) When controlling for focal paper’s novelty (panel A), interdisciplinarity (panel B), reference age (panel C), reference impact $C_{5}$ (panel D), and reference disruption (panel E), respectively, both the percentile (blue curves) and positive ratio (green curves) of disruption continue to increase with KI across all levels of above covariates.
Figure S25: $|$ The association between disruption and KI is robust when controlling for team compositions in Web of Science dataset. (A-C) When controlling for team size (panel A), geographic distance (panel B), and collaboration freshness (panel C) separately, both the percentile (blue curves) and positive ratio (green curves) of disruption continue to increase with KI across all levels of above team compositions. Additionally, papers produced by small, onsite, and fresh teams tend to exhibit slightly higher disruption. Considering the lack of coordinate information in Web of Science dataset, we categorize the team distance of a given focal paper by hierarchical categorization of the affiliations: 0 (Sub-organization / Department level): 11: Organization / University level; 22: City level; 33: State / Province level; 44: National level; 55: International level.
Figure S26: $|$ Papers with higher KI are underrepresented among cited references and receiving delayed citations in Web of Science dataset. (A-F) For papers published in the year $T$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline.
Table S11: $|$ Normalized regression models of disruption and the influencing factors in Web of Science dataset.
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 Model 9 $R^{2}$%
KI 0.4204*** 0.4203*** 0.4189*** 0.4081*** 0.4184*** 0.4203*** 0.4186*** 0.4204*** 0.4036*** 71.16
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
Novelty 0.0063*** 0.0020*** 0.04
@@
Interdisciplinarity 0.0164*** 0.0168*** 0.74
(0.000) (0.000)
Reference count -0.0364*** -0.0342*** 9.56
(0.000) (0.000)
Reference age 0.0538*** 0.0508*** 0.85
(0.000) (0.000)
Team size -0.0061*** -0.0034*** 0.07
(0.000) (0.000)
Team distance -0.0330*** -0.0308*** 2.77
(0.000) (0.000)
Team freshness 0.0002 0.0002 0.35
(0.000) (0.000)
Discipline fixed effects Yes Yes Yes Yes Yes Yes Yes Yes Yes 7.30
Year fixed effects Yes Yes Yes Yes Yes Yes Yes Yes Yes 7.16
N 12145196 12145196 12145196 12145196 12145196 12145196 12145196 12145196 12145196
$R^{2}$ 0.221 0.221 0.221 0.222 0.224 0.221 0.222 0.221 0.226
  • $\dagger$
    We perform normalized ordinary-least-squares (OLS) regression analyses to examine the effect of KI on disruption in Web of Science dataset, while controlling for various covariates. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S12: $|$ Normalized mediation regression on team size in Web of Science dataset.
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$) ||ACME / TE||
Team size KI -0.0043*** -0.0064*** -0.0107*** 0.4029
@@
Team size Novelty -0.0001*** -0.0106*** -0.0107*** 0.0139
@@
Team size Interdisciplinarity 0.0002*** -0.0109*** -0.0107*** 0.0149
@@
Team size Reference count -0.0006*** -0.0101*** -0.0107*** 0.0577
@@
Team size Reference age -0.0015*** -0.0093*** -0.0107*** 0.1359
@@
Team size Team distance -0.0037*** -0.0070*** -0.0107*** 0.3443
@@
Team size Team freshness -0.0000*** -0.0107*** -0.0107*** 0.0000
@@
  • $\dagger$
    We perform normalized mediation analysis to assess the effect of team size on disruption in Web of Science dataset. The candidate mediating variables include: KI, novelty, interdisciplinarity, reference count, reference age, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team size, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S13: $|$ Normalized mediation regression on team distance in Web of Science dataset.
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$) ||ACME / TE||
Team distance KI -0.0254*** -0.0298*** -0.0553*** 0.4602
@@
Team distance Novelty -0.0000*** -0.0552*** -0.0553*** 0.0008
@@
Team distance Interdisciplinarity 0.0023*** -0.0576*** -0.0553*** 0.0422
@@
Team distance Reference count -0.0085*** -0.0468*** -0.0553*** 0.1536
@@
Team distance Reference age -0.0021*** -0.0532*** -0.0553*** 0.0379
@@
Team distance Team size -0.0005*** -0.0548*** -0.0553*** 0.0086
@@
Team distance Team freshness -0.0000*** -0.0553*** -0.0553*** 0.0000
@@
  • $\dagger$
    We perform normalized mediation analysis to assess the effect of team distance on disruption in Web of Science dataset. The candidate mediating variables include: KI, novelty, interdisciplinarity, reference count, reference age, team size, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team distance, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S14: $|$ Normalized mediation regression on team freshness in Web of Science dataset.
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$) ||ACME / TE||
Team freshness KI 0.0003*** 0.0006*** 0.0008*** 0.3044
@@
Team freshness Novelty 0.0000*** 0.0008*** 0.0008*** 0.0014
@@
Team freshness Interdisciplinarity -0.0000*** 0.0009*** 0.0008*** 0.0401
@@
Team freshness Reference count 0.0002*** 0.0006*** 0.0008*** 0.2572
@@
Team freshness Reference age -0.0000 0.0008*** 0.0008*** 0.0010
@@
Team freshness Team size 0.0000*** 0.0008*** 0.0008*** 0.0016
@@
Team freshness Team distance 0.0001*** 0.0007*** 0.0008*** 0.1398
@@
    We perform normalized mediation analysis to assess the effect of team freshness on disruption in Web of Science dataset. The candidate mediating variables include: KI, novelty, interdisciplinarity, reference count, reference age, team size, and team distance, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team freshness, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table S15: $|$ Normalized confounding regression on paper impact in Web of Science dataset.
Independent variable Confounder Unadjusted ($\alpha_{1}$) Adjusted ($\delta_{1}$) RCE ($|\alpha_{1}-\delta_{1}|/|\alpha_{1}|$)
Citation percentile KI -0.1302*** -0.0276*** 0.7878
(0.0006) (0.0006)
Citation percentile Novelty -0.1302*** -0.1305*** 0.0022
(0.0006) (0.0006)
Citation percentile Interdisciplinarity -0.1302*** -0.1328*** 0.0202
(0.0006) (0.0006)
Citation percentile Reference count -0.1302*** -0.0872*** 0.3306
(0.0006) (0.0009)
Citation percentile Reference age -0.1302*** -0.1217*** 0.0649
(0.0006) (0.0006)
Citation percentile Team size -0.1302*** -0.1300*** 0.0016
(0.0006) (0.0006)
Citation percentile Team distance -0.1302*** -0.1257*** 0.0345
(0.0006) (0.0006)
Citation percentile Team freshness -0.1302*** -0.1302*** 0.0000
(0.0006) (0.0006)
  • $\dagger$